### v1

@@ 2 Problem Formulation
The significance of this metric is evident. Based on the definition (12), the malfunctioning system (5) will use at most $r_{A}(t_{f},R)$ more energy than the nominal system (1) to achieve finite-time reachability, for a given final time $t_{f}$ and $x_{0}$ no farther than $R$ from $x_{tg}$. The constraint $\|x_{0}-x_{tg}\|_{2}\leq R$ is required to ensure that the value of $r_{A}(t_{f},R)$ is finite.
Remark 1.
@@
In the nonlinear systems (1) and (5), deriving a closed-form expression for the control that achieves $x(t_{f})=x_{tg}$ is generally intractable (see, e.g., [35, 36]). Similarly, closed-form expressions for the minimum energies (8) and (9) and subsequently the total (10) and worst-case total (11) energies are generally not derivable [37]. To address this, we derive a condition on the mean value of the control in both the nominal and malfunctioning cases in Section 3. This condition achieves our first objective in Problem 1, and is used to approximate the minimum energies and worst-case total energy, and these approximations are used to derive a bound on the energetic resilience metric in Section 4.
3 Approximations for Nominal, Malfunctioning and Worst-case Total Energies
In this section, we derive approximations for the nominal, malfunctioning and worst-case total energies defined in Section 2. This approximation uses a characterization of the mean value of the control signal in the interval $[0,t_{f}]$, which is derived from the system equations (1) and (5).
3.1 Nominal and Malfunctioning Energies
@@
$x(t)=x_{0}+\int_{0}^{t}f(x(\tau))\mathrm{d}\tau+\int_{0}^{t}g(x(\tau))u(\tau)\mathrm{d}\tau.$ (13)
Let $\tilde{x}\coloneqq x_{0}-x_{tg}$, $B=g(x_{0})$ and $f_{0}=f(x_{0})$. If $x(t_{f})=x_{tg}$,
$\tilde{x}+t_{f}B\overline{u}=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t,$ (14)
where we use the property (3) for $g$, and $\overline{u}=\frac{1}{t_{f}}\int_{0}^{t_{f}}u(t)\mathrm{d}t$ denotes the mean value of $u$ in the interval $[0,t_{f}]$. Let $v\equiv v(t_{f},u)\coloneqq-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$. If $v$ was a fixed quantity with respect to $u$ and hence $\overline{u}$, (14) is linear in $\overline{u}$ and we can write the following least-norm solution [38]:
$\overline{u}^{LS}=\frac{1}{t_{f}}B^{\dagger}(v-\tilde{x}),$ (15)
@@
The quantity $v=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$ is bounded as
$\|v\|_{\infty}\leq\overline{v}\coloneqq\frac{1}{D_{f}}\left[c(e^{t_{f}D_{f}}-1)-t_{f}D_{f}\|B\|_{\infty}\right]$ (16)
where $c=\|f_{0}\|_{\infty}+\|B\|_{\infty}+D_{g}$.
@@
for each $i=1,\ldots,m+p$, where $[\cdot]_{i}$ denotes the $i$-th component of the argument. Each of the above conditions is linear in $v$, and thus can be checked efficiently on the vertices of the hypercube $\|v\|_{\infty}\leq\overline{v}$ [39].
@@
For a piecewise continuous function $z:[0,t_{f}]\to\mathbb{R}^{n_{z}}$ with a given mean value $\overline{z}\in\mathbb{R}^{n_{z}}$, the energy of $z$ is related to its mean value by
@@
Since the squared norm $\|\cdot\|_{2}^{2}$ is a convex function, we have $\|\overline{z}\|_{2}^{2}=\left\|\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)\mathrm{d}t\right\|_{2}^{2}\leq\frac{1}{t_{f}}\int_{0}^{t_{f}}\|z(t)\|_{2}^{2}\mathrm{d}t,$ using Jensen’s inequality [32]. From this, the result follows. ∎
Any control signal $u$ that achieves finite-time reachability must satisfy (14). Subsequently, its mean value $\overline{u}$ must satisfy (15). Thus, the optimal control signal with minimum energy $E_{N}^{*}(x_{0},x_{tg},t_{f})$ that achieves finite-time reachability must have mean value $\overline{u}^{LS}$, where we recall that $\|v\|_{\infty}\leq\overline{v}$. Using Lemma 1, $E_{N}^{*}(x_{0},x_{tg},t_{f})\geq t_{f}\|\overline{u}^{LS}\|_{2}^{2}$. We now make the central approximation in this article, namely, that
$E_{N}^{*}(x_{0},x_{tg},t_{f})\approx t_{f}\|\overline{u}^{LS}\|_{2}^{2}.$ (19)
Note that this approximation is exact when the optimal control is constant, i.e., $u^{*}(t)=\overline{u}^{LS}$ for all $t\in[0,t_{f}]$. In particular, this approximation is exact in the case of linear, driftless systems where $D_{f}=D_{g}=0$ and $f_{0}=0$, as shown in [34] where the optimal control $u^{*}(t)=\overline{u}_{LS}$ for all $t\in[0,t_{f}]$. We also use such an approximation later in the malfunctioning case. These approximations are used to circumvent the difficulties in obtaining closed-form expressions for optimal control signals and the corresponding minimum energies [37]. Under the assumption that this approximation is exact, we obtain
@@
$x(t)=x_{0}+\int_{0}^{t}f(x(\tau))\mathrm{d}\tau+\int_{0}^{t}g_{c}(x(\tau))u_{c}(\tau)\mathrm{d}\tau+\int_{0}^{t}g_{uc}(x(\tau))u_{uc}(\tau)\mathrm{d}\tau.$ (21)
Let $B_{c}\coloneqq g_{c}(x_{0})$ and $B_{uc}\coloneqq g_{uc}(x_{0})$. Then, under finite-time reachability with $x(t_{f})=x_{tg}$,
$\tilde{x}+t_{f}B_{c}\overline{u}_{c}+t_{f}B_{uc}\overline{u}_{uc}=-\int_{0}^{t}f(x(\tau))\mathrm{d}\tau-\int_{0}^{t_{f}}d_{g}(x(t))u(t)\mathrm{d}t$ (22)
where $u(t)=\left[u_{c}^{\operatorname*{\mathsf{T}}}(t),u_{uc}^{\operatorname*{\mathsf{T}}}(t)\right]^{\operatorname*{\mathsf{T}}}$, and we define the mean values $\overline{u}_{c}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u_{c}(t)\mathrm{d}t$ and $\overline{u}_{uc}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u_{uc}(t)\mathrm{d}t$. Note that the right-hand side equals $v$ as defined earlier. For a fixed $v$, the above equation is linear in $\overline{u}_{c}$, and similar to the nominal case, we have the least-norm solution
@@
where $\|v\|_{\infty}\leq\overline{v}$ from (16). We remind the reader that $v$ is not a fixed quantity, but the norm of $\overline{u}_{c}^{LS}$ can be used to approximate the malfunctioning energy $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ in (9) as $v$ is bounded. First, we investigate conditions under which the constraint $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ is satisfied using (23) and the bound $\overline{v}$.
@@
Finally, we consider the malfunctioning energy $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ in (9) and note that obtaining an exact expression is intractable [37] for the general nonlinear system (5). We thus use (23) to approximate $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ using Lemma 1. From (18), we have $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\geq t_{f}\|\overline{u}_{c}^{LS}\|_{2}^{2}$, as the mean of the optimal control signal in (9) must satisfy (23) to achieve finite-time reachability. Then, similar to the approximation for the nominal energy (19), we write the approximate malfunctioning energy
@@
where we recall that $\overline{u}_{c}^{LS}$ depends on $v(t_{f},u)$, which is bounded as in (16). As before, this approximation is exact when the optimal control is constant, i.e., $u_{c}^{*}(t)=\overline{u}_{c}^{LS}$ for all $t\in[0,t_{f}]$, which is true in the case of linear, driftless systems as shown in [34]. Under the assumption that this approximation is exact, we obtain
@@
Equations (20) and (26) provide the approximations for the nominal and malfunctioning energies and are obtained using Proposition 1 and Lemma 1. The conditions (17) and (24) are used to check whether the constraints $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ are satisfied for a given $t_{f}$. In the next subsection, we use these results to obtain an approximate upper bound on the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ defined in (11).
3.2 Worst-case Total Energy

### v3

@@ 2. Problem Formulation
The significance of this metric is evident. Based on the definition (12), the malfunctioning system (5) will use at most $r_{A}(t_{f},R)$ more energy than the nominal system (1) to achieve fixed-time reachability, for a given final time $t_{f}$ and $x_{0}$ no farther than $R$ from $x_{tg}$. The constraint $\|x_{0}-x_{tg}\|_{2}\leq R$ is required to ensure that the value of $r_{A}(t_{f},R)$ is finite.
We note that an alternative metric for energetic resilience was considered in [27]. However, the treatment in [27] considered only linear driftless systems, and the corresponding metric is not meaningful for general nonlinear systems due to differences in system responses as discussed in Section 4. In contrast, the metric (12) we consider in this paper can be extended to nonlinear systems. In the next section, we also begin by considering linear driftless systems and deriving a bound on the metric (12). The results in this section are used as a basis for addressing resilience of general nonlinear systems in Section 4 and beyond.
3. Resilience of Linear Driftless Systems
In this section, we consider the special case of linear driftless systems, i.e., $f(x(t))\equiv 0$ and $g(x(t))\equiv B$ for all $x(t)$ in (1). We first review the expressions for the nominal (8), malfunctioning (9) and worst-case total (11) energies for linear driftless systems derived in [27]. Based on these, we briefly consider the problem of designing the optimal final time to minimize malfunctioning energy. We then address the central problem of bounding the metric (12), in the case of losing authority over one actuator. In the case of linear driftless systems, we have $f(x(t))\equiv 0$ and $g(x(t))\equiv B$ so that $D_{f}=D_{g}=0$ and the system (1) reduces to
$\dot{x}(t)=Bu(t),~x(0)=x_{0}.$ (13)
Thus, under a malfunction of the form described in Section 2, the malfunctioning system (5) is written as
$\dot{x}(t)=B_{c}u_{c}(t)+B_{uc}u_{uc}(t).$ (14)
Under the assumption that the nominal dynamics (13) are controllable, it was shown that a control input $u^{*}$ achieves the task if and only if
$u^{*}(t)=\overline{u}^{LS}=-\frac{1}{t_{f}}B^{\dagger}\tilde{x}~\text{ for all }~t\in[0,t_{f}],$ (15)
where $\overline{u}^{LS}$ denotes the mean value of the control with the minimum energy. Note that this control is constant in the interval $0,t_{f}]$. Further, the constraint $u\in\operatorname*{\mathcal{U}}$, i.e., $\|u(t)\|_{\infty}\leq 1$ is satisfied if and only if
$t_{f}\geq\|B^{\dagger}\tilde{x}\|_{\infty}$. (16)
Finally, the nominal energy was written as
$E_{N}^{*}(x_{0},x_{tg},t_{f})=\int_{0}^{t_{f}}u^{*T}(t)u^{*}(t)\mathrm{d}t=\frac{1}{t_{f}}\|B^{\dagger}\tilde{x}\|_{2}^{2}.$ (17)
Using similar arguments, the malfunctioning energy (9) was derived in [27] as
$E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(\tilde{x}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}$ (18)
under conditions similar to (16) on the final time $t_{f}$. From this, the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ was shown to be bounded from above as follows [27, Proposition 1]:
$\overline{E}_{T}(x_{0},x_{tg},t_{f})$ $\leq\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+t_{f}\left(\sum_{i=1}^{p}{}_{i}\|V\|_{1}^{2}+p\right)+2\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\|_{1}.$ (19)
In the particular case of losing control authority over one actuator, i.e., $p=1$, it was also shown that the following closed-form expression can be obtained for the worst-case total energy:
$\overline{E}_{T}(x_{0},x_{tg},t_{f})$ $=\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)+2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|.$ (20)
3.1. Final Time Design
In the context of the malfunctioning energy (18), it is useful to investigate the optimal final time $t_{f}^{*}$ that minimizes this energy for a given uncontrolled input $u_{uc}$. This is useful in designing control signals for tasks with a variable completion time and constraints on the total available control energy. In other words, we wish to solve the following problem:
$t_{f}^{*}$ $=\operatorname*{arg\,min}_{t_{f}}E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$
$=\operatorname*{arg\,min}_{t_{f}}\frac{1}{t_{f}}\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}.$ (21)
Using the first-order stationary condition $\frac{\partial}{\partial t_{f}}\left(E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\right)=0$, we have
$t_{f}\frac{\partial}{\partial t_{f}}\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2}=\left\|B_{c}^{\dagger}\left(x_{0}+t_{f}B_{uc}\overline{u}_{uc}\right)\right\|_{2}^{2},$
with $t_{f}\neq 0$. The above equality simplifies to the following quadratic equation in $t_{f}$:
$t_{f}^{2}\overline{u}_{uc}^{T}B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}B_{uc}\overline{u}_{uc}=x_{0}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}.$
Solving for $t_{f}>0$, we have:
$t_{f}^{*}=\frac{\|B_{c}^{\dagger}x_{0}\|_{2}}{\|B_{c}^{\dagger}B_{uc}\overline{u}_{uc}\|_{2}}.$ (22)
It is then straightforward to show that
$\frac{\partial^{2}}{\partial t_{f}^{2}}\left(E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\right)=\frac{2\|B_{c}^{\dagger}x_{0}\|_{2}^{2}}{t_{f}^{3}}~\text{ and thus}$~
$\frac{\partial^{2}}{\partial t_{f}^{2}}\left(E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\right)\Bigg|_{t_{f}=t_{f}^{*}}=\frac{2\|B_{c}^{\dagger}B_{uc}\overline{u}_{uc}\|_{2}^{3}}{\|B_{c}^{\dagger}x_{0}\|_{2}}>0.$
Hence $t_{f}^{*}$ in (22) is the unique minimum for the problem in (3.1). The result (22) thus provides the final time that minimizes the malfunctioning energy for a given uncontrolled input $\overline{u}_{uc}$.
3.2. Resilience of Linear Driftless Systems to the Loss of One Actuator
We now return to the expressions (17)-(20) and derive a bound on the energetic resilience metric (12) under loss of control authority over one actuator. We note that a bound on $r_{A}(t_{f},R)$ can be derived when control authority is lost over $p>1$ actuators using (19). However, this bound will be more conservative as (19) is only an upper bound an not an exact expression as in (20).
Proposition 1.
For the case when $p=1$, the additive energetic resilience metric $r_{A}(t_{f},R)$ is bounded from above as follows:
$r_{A}(t_{f},R)$ $\leq\frac{R^{2}}{t_{f}}{}_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)+2R\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right).$ (23)
Proof.
Using (17), (20) and (12),
$r_{A}(t_{f},R)$ $=\sup_{\|x_{0}\|_{2}\leq R}\Biggl\{\frac{1}{t_{f}}\left\|B_{c}^{\dagger}x_{0}\right\|_{2}^{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)+2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|-\frac{1}{t_{f}}\left\|B^{\dagger}x_{0}\right\|_{2}^{2}\Biggr\}$
$\leq t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right)+2R\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}+\sup_{\|x_{0}\|_{2}\leq R}\frac{1}{t_{f}}x_{0}^{T}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)x_{0},$ (24)
where the inequality follows from the Cauchy-Schwarz inequality applied to the term $2\left|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}x_{0}\right|$. Note that $B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}$ is symmetric, and hence using the Rayleigh inequality and (24),
$r_{A}(t_{f},R)$ $\leq\frac{R^{2}}{t_{f}}{}_{\max}\left(B_{c}^{\dagger T}B_{c}^{\dagger}-B^{\dagger T}B^{\dagger}\right)+2R\|B_{uc}^{T}B_{c}^{\dagger T}B_{c}^{\dagger}\|_{2}+t_{f}\left(\|B_{uc}\|_{2}^{2}+1\right),$
concluding the proof. ∎
In the following section, we begin considering the general nonlinear systems (1) and (5), deriving approximations for the energies defined in Section 2. These approximations rely on the difference in responses between the general nonlinear system (1) and the linear driftless system (13).
4. Approximations for Nominal, Malfunctioning and Worst-case Total Energies
@@
We now consider the general nonlinear systems (1) and (5), and derive approximations for the nominal, malfunctioning and worst-case total energies defined in Section 2. We note that for these nonlinear systems, deriving a closed-form expression for the control that achieves $x(t_{f})=x_{tg}$ is generally intractable (see, e.g., [39, 40]). Similarly, closed-form expressions for the minimum energies (8) and (9) and subsequently the total (10) and worst-case total (11) energies can generally not be derived [41] in the nonlinear case. To address this, we derive a condition on the mean value of the control in the interval $0,t_{f}]$, in both the nominal and malfunctioning cases. This condition is the solution to Problem 1, and is used to approximate the minimum energies and worst-case total energy. These approximations are used to derive a bound on the energetic resilience metric in Section 5.
4.1. Nominal and Malfunctioning Energies
@@
[…]
Let $\tilde{x}\coloneqq x_{0}-x_{tg}$, $B=g(x_{0})$ and $f_{0}=f(x_{0})$. We assume $B$ has full row rank, similar to the assumption made in Section 3. If $x(t_{f})=x_{tg}$,
$\tilde{x}+t_{f}B\overline{u}=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t,$ (26)
where we use the property (3) for $g$, and $\overline{u}=\frac{1}{t_{f}}\int_{0}^{t_{f}}u(t)\mathrm{d}t$ denotes the mean value of $u$ in the interval $0,t_{f}]$. Let $v\equiv v(t_{f},u)\coloneqq-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t$. If $v$ was a fixed quantity with respect to $u$ and hence $\overline{u}$, (26) is linear in $\overline{u}$ and we can write the following least-norm solution [42]:
[…]
@@
The quantity $v=-\int_{0}^{t_{f}}f(x(t))\mathrm{d}t-\int_{0}^{t_{f}}d_{g}(x(t),x_{0})u(t)\mathrm{d}t$ is bounded as
$\|v\|_{\infty}\leq\overline{v}\coloneqq\frac{1}{D_{f}+D_{g}}\left[c\left(e^{t_{f}(D_{f}+D_{g})}-1\right)-t_{f}(D_{f}+D_{g})\|B\|_{\infty}\right]$ (28)
where $c=\|f_{0}\|_{\infty}+\|B\|_{\infty}$.
@@
It is clear that $v(t_{f},u)$ denotes the difference between system responses under the driftless dynamics (13) and the general nonlinear dynamics (1) for a given control input $u\in\operatorname*{\mathcal{U}}$. Proposition 2 thus shows that this difference between system responses is bounded and quantified by $\overline{v}$, where $\overline{v}$ is a function of the system dynamics, initial condition and final time. In particular, when $f(x(t))\equiv 0$ and $g(x(t))\equiv B$, the dynamics in (1) reduce to (13) and we obtain $v(t_{f},u)=0$, $\overline{v}=0$ since $D_{f}=D_{g}=0$ and $f_{0}=0$. Next, let $D_{S}\coloneqq D_{f}+D_{g}$. In investigating how the difference in system responses between (1) and (13) changes as $D_{f}$ and $D_{g}$ increase, we note that
$\frac{\partial\overline{v}}{\partial D_{S}}=\frac{c[(t_{f}D_{S}-1)e^{t_{f}D_{S}}+1]}{D_{S}^{2}}\geq 0,$
i.e., $\overline{v}$ is a non-decreasing function of $D_{S}$. Thus, the difference in system responses increases as $D_{S}$ increases. The facts discussed above are used to motivate the approximations for different energies derived later in this section.
@@
for each $i=1,\ldots,m+p$, where $\cdot]_{i}$ denotes the $i$-th component of the argument. Each of the above conditions is linear in $v$, and thus can be checked efficiently on the vertices of the hypercube $\|v\|_{\infty}\leq\overline{v}$ [43].
@@
For a piecewise continuous function $z:[0,t_{f}]\to\mdmathbb{R}^{n_{z}}$ with a given mean value $\overline{z}\in\mdmathbb{R}^{n_{z}}$, the energy of $z$ is related to its mean value by
@@
The squared norm $\|\cdot\|_{2}^{2}$ is a convex function, Using Jensen’s inequality [37], we have $\|\overline{z}\|_{2}^{2}=\left\|\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)\mathrm{d}t\right\|_{2}^{2}\leq\frac{1}{t_{f}}\int_{0}^{t_{f}}\|z(t)\|_{2}^{2}\mathrm{d}t$. From this, the result follows. ∎
Any control signal $u$ that achieves fixed-time reachability must satisfy (26), where $\overline{u}$ is its mean value. On the basis of Lemma 1, we use the right-hand side of (30) to approximate the energy in $u$. Further, based on [27, Lemma 1], we are interested in the mean value with minimum-norm that solves (26). This quantity is given by $\overline{u}^{LS}$ as defined in (27) for a particular $v$. We now make the central approximation in this article, namely, that
$E_{N}^{*}(x_{0},x_{tg},t_{f})\approx t_{f}\|\overline{u}^{LS}\|_{2}^{2},$ (31)
thus approximating (30) with equality using the minimum-norm mean $\overline{u}^{LS}$. We remark that the conservatism of this approximation can be measured by comparing it to the energy of any control signal $u\in\operatorname*{\mathcal{U}}$ that achieves the objective $x(t_{f})=x_{tg}$.
Note that the approximation (31) is exact when the optimal control is constant, i.e., $u^{*}(t)=\overline{u}^{LS}$ for all $t\in[0,t_{f}]$. In particular, this approximation is exact in the case of linear driftless systems where $D_{f}=D_{g}=0$ and $f_{0}=0$, as shown in [27] where the optimal control $u^{*}(t)=\overline{u}_{LS}$ for all $t\in[0,t_{f}]$. In this case, the quantities $v(t_{f},u)$ and its bound $\overline{v}$ reduce to zero. For general nonlinear systems, this approximation is more accurate for smaller values of $D_{f}$ and $D_{g}$, as such settings result in small values of $\overline{v}$. Here, we recall that $\overline{v}$ quantifies the difference in system responses between (1) and (13), and $\overline{v}$ is a non-decreasing function of $D_{S}$. We also use an approximation similar to (31) later in the malfunctioning case. These approximations are used to circumvent the difficulties in obtaining closed-form expressions for optimal control signals and the corresponding minimum energies [41]. Under the assumption that this approximation is exact, we obtain the nominal energy
@@
[…]
Let $B_{c}\coloneqq g_{c}(x_{0})$ and $B_{uc}\coloneqq g_{uc}(x_{0})$. Then, under fixed-time reachability with $x(t_{f})=x_{tg}$,
[…]
where $u(t)=\left[u_{c}^{\operatorname*{\mathsf{T}}}(t),u_{uc}^{\operatorname*{\mathsf{T}}}(t)\right]^{\operatorname*{\mathsf{T}}}$, $\overline{u}_{c}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u_{c}(t)\mathrm{d}t$ and $\overline{u}_{uc}\coloneqq\frac{1}{t_{f}}\int_{0}^{t_{f}}u_{uc}(t)\mathrm{d}t$ are the mean values of the controlled and uncontrolled inputs. As with the nominal case and similar to [27], we assume $B_{c}$ has full row rank. Note that the right-hand side equals $v$ as defined earlier. For a fixed $v$, the above equation is linear in $\overline{u}_{c}$, and similar to the nominal case, we have the least-norm solution
@@
where $\|v\|_{\infty}\leq\overline{v}$ from (28). We note here that (27) and (35) provide a characterization of the controls ensuring $x_{tg}$ is reachable and resiliently reachable, thus addressing Problem 1. We also remind the reader that $v$ is not a fixed quantity, but the norm of $\overline{u}_{c}^{LS}$ can be used to approximate the malfunctioning energy $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ in (9) as $v$ is bounded. First, we investigate conditions under which the constraint $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ is satisfied using (35) and the bound $\overline{v}$.
@@
Finally, we consider the malfunctioning energy $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ in (9) and note that obtaining an exact expression is intractable [41] for the general nonlinear system (5). We thus use (35) to approximate $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})$ using Lemma 1. From (30), we have $E_{M}^{*}(x_{0},x_{tg},t_{f},u_{uc})\geq t_{f}\|\overline{u}_{c}^{LS}\|_{2}^{2}$, where we choose the mean of the optimal control signal to satisfy (35), motivated by [27, Lemma 1]. Then, similar to the approximation for the nominal energy (31), we write the approximate malfunctioning energy
@@
where we recall that $\overline{u}_{c}^{LS}$ depends on $v(t_{f},u)$, which is bounded as in (28). As before, this approximation is exact when the optimal control is constant, i.e., $u_{c}^{*}(t)=\overline{u}_{c}^{LS}$ for all $t\in[0,t_{f}]$, which is true in the case of linear driftless systems as shown in [27]. Under the assumption that this approximation is exact, we obtain
@@
Equations (32) and (38) provide the approximations for the nominal and malfunctioning energies and are obtained using Proposition 2 and Lemma 1. In Appendix B, we list all approximate and non-approximate quantities in this paper for clarity. The conditions (29) and (36) are used to check whether the constraints $u\in\operatorname*{\mathcal{U}}$ and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ are satisfied for a given $t_{f}$. In the next subsection, we use these results to obtain an approximate upper bound on the worst-case total energy $\overline{E}_{T}(x_{0},x_{tg},t_{f})$ defined in (11).
4.2. Worst-case Total Energy